OpenStudy (anonymous):

Solve. 13-1/5(45w+30)=-10w-9

OpenStudy (anonymous):

13-8w-6=-10w-9

OpenStudy (anonymous):

13-6+9=-10w+8w

OpenStudy (anonymous):

16=-2w

OpenStudy (anonymous):

-2w/2=16/-2

OpenStudy (anonymous):

w=-8